Google Chrome Secure I gnment ProblemID 76845612 Exercise 38.24 Part A horizontal beam of laser light of wavelength 613 nmpasses has o 0620 mm The inensity of What is the minimum uncertainty inthe vertcal component ofthe momentum ofeach in the beam later the photon has passed sat? Express your answer with the appropriate unts nroughte the light is measured on a vertcal screen thatis 300 mtom he kem AP, 851-10-11 Mr. Give up Part B use he result orpartAtoestmate the wdth of central dtadon maodmum matis observed on the soren Express your answer to two signincant moures and clude the appropriate unts. Value Incorrect: Try Again: 4 attempts remaining a OAsk me anything
Answer #1

del px = h/ lamda = 6.626 * 10^-34/613 * 10^-9 = 0.0108 * 10^-25

theta = py/px = 8.51 * 10^-31/1.08 * 10^-27=7.87 * 10^-4 rad

the width is

W= 3.00 ( 7.87 * 10^-4 rad) = 2.363 mm